OpenStudy (anonymous):

the lines 4x + 3y = 2 and 8x + 6y = 3 are ________?

OpenStudy (anonymous):

might want to get it in this form \[y=mx+b\]

OpenStudy (anonymous):

Maybe you might spot something then.

OpenStudy (anonymous):

sniping?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ax+by=c lx + my =n They are parallel when \(\large \frac a b = \frac l m \neq \frac c n \)

OpenStudy (anonymous):

ok